OpenStudy (anonymous):

What is the simplified form of the expression sqrt 16 c^8-sqrt c^2+2 sqrt^2+sqrt 4^2

OpenStudy (anonymous):

u talking about this: \(\large \sqrt{16c^8}-\sqrt{c^2}+2\sqrt2+\sqrt{4^2} \)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\(\large \sqrt{16c^8}-\sqrt{c^2}+2\sqrt2+\sqrt{4^2} \) \(\large 4c^4-c+2\sqrt2+4 \) that's as much as you can do with that... can you double check that third term where you have "2sqrt^2" .... is that right?

OpenStudy (anonymous):

or better yet, use the draw to write out the expression you want...

OpenStudy (anonymous):